The length of a sign is 4 inches more than the width. The area is 45 square inches. Identify the length and the width of the sign.
Thank you

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

I am going to guess that this sign is rectangular.
Length(l) = w + 4
Width(w) = w
Length * Width = Area
(w + 4)(w) = 45
w^2 + 4w = 45
w^2 + 4w - 45 = 0
(w + 9)(w - 5) = 0
w+=+-9 or w+=+5
Since the width can not be negative, your width is: w+=+5
Plug:
Length = w + 4
Length = 5 + 4
Length = 9

Answer by checkley71(8403)

(4+X)*X=45 OR 4X+X~2-45=0 OR (X-5)(X+9)=0 OR X=5 & X=-9